Can you help me solve this problem please!!!
3/4-(a-2)-1/3(5-2a)=-2

Answer by Alan3354(69443)

3/4-(a-2)-1/3(5-2a)=-2
3/4 -a+2 - 1/3(5-2a)=-2
Multiply by 12
9 - 12a + 24 - 4(5-2a) = -24
33 - 12a - 20 + 8a = -24
13 - 4a = -24
-4a = -37
a = 37/4